the perimeter of a rectangle is 50cm if one side is 3x-8 and the other side is x+9 what does x equal
Answer by waynest(281)

2(3x - 8) + 2(x + 9) = 50
6x - 16 + 2x + 18 = 50
8x + 2 = 50
8x + 2 - 2 = 50 - 2
8x = 48
8x/8 = 48/8
x = 6
check:
2(3(6) - 8) + 2(6 + 9) = 50
2(18 - 8) + 2(15) = 50
20 + 30 = 50
50 = 50